OpenStudy (anonymous):

-4x + 7y = 1 25 = 2x + 5y find dx and d whomever helps me will receive a medal

OpenStudy (anonymous):

-4x + 7y = 1 25 - 2x - 5y = 0 -4x + 7y = 1 -2x - 5y = _25 thats what i got so far

OpenStudy (anonymous):

-4x + 7y = 1 _2, * _2 x + -2 * -5y + -2 (25)

OpenStudy (anonymous):

i didnt slove my own question